# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ Appendix G Technical Lemmas for Near-Constant Weight Designs
$\mathbb{P}\big{(}U=C\Delta\big{)}=\exp\big{(}-\Delta\big{[}D(C\|e^{-\nu})+o(1)\big{]}\big{)}.$ (309)
@@
$\mathbb{P}(U=C\Delta)=\frac{\binom{k_{\nu}\Delta}{C\Delta}\binom{(k_{0}-k_{\nu})\Delta}{(1-C)\Delta}}{\binom{k_{0}\Delta}{\Delta}}.$ (310)
Given positive integers $m$ and $\alpha m$ for some $\alpha\in(0,1)$, we have $\frac{\sqrt{\pi}}{2}G\leq\binom{m}{\alpha m}\leq G$ where $G=\frac{\exp(mH_{2}(\alpha))}{\sqrt{2\pi m\alpha(1-\alpha)}}$ (see Lemma 2), which gives
@@
Substituting this into (310), we obtain
$\mathbb{P}\big{(}U=C\Delta\big{)}$ $=\Theta(1)\frac{\frac{\exp(k_{\nu}\Delta H_{2}(\frac{C}{k_{\nu}}))}{\sqrt{C\Delta(1-\frac{C}{k_{\nu}})}}\cdot\frac{\exp((k_{0}-k_{\nu})\Delta H_{2}(\frac{1-C}{k_{0}-k_{\nu}}))}{\sqrt{(1-C)\Delta(1-\frac{1-C}{k_{0}-k_{\nu}})}}}{\frac{\exp(k_{0}\Delta H_{2}(\frac{1}{k_{0}}))}{\sqrt{\Delta(1-\frac{1}{k_{0}})}}}$ (311)
$=\frac{\Theta(1)}{\sqrt{C(1-C)\Delta}}\exp\Big{(}\Delta\Big{[}\underbrace{k_{\nu}H_{2}\Big{(}\frac{C}{k_{\nu}}\Big{)}+(k_{0}-k_{\nu})H_{2}\Big{(}\frac{1-C}{k_{0}-k_{\nu}}\Big{)}-k_{0}H_{2}\Big{(}\frac{1}{k_{0}}\Big{)}}_{:=\mathcal{T}}\Big{]}\Big{)},$ (312)
where (312) follows from $1-\frac{C}{k_{\nu}}$, $1-\frac{1-C}{k_{0}-k_{\nu}}$, and $1-\frac{1}{k_{0}}$ all behaving as $\Theta(1)$. Now we seek to further simplify $\mathcal{T}$. Note that $\frac{C}{k_{\nu}}$, $\frac{1-C}{k_{0}-k_{\nu}}$, and $\frac{1}{k_{0}}$ all scale as $o(1)$, and it holds for any $t=o(1)$ that
$H_{2}(t)$ $=-t\log t-(1-t)\log(1-t)=-t(\log t-1+o(1)).$ (313)
@@
$\mathcal{T}$ $=k_{\nu}\Big{(}-\frac{C}{k_{\nu}}\Big{[}\log\Big{(}\frac{C}{k_{\nu}}\Big{)}-1+o(1)\Big{]}\Big{)}-(k_{0}-k_{\nu})\cdot\frac{1-C}{k_{0}-k_{\nu}}\cdot\Big{[}\log\Big{(}\frac{1-C}{k_{0}-k_{\nu}}\Big{)}-1+o(1)\Big{]}$ (314)
$\quad\quad\quad\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +k_{0}\cdot\frac{1}{k_{0}}\cdot\Big{[}\log\Big{(}\frac{1}{k_{0}}\Big{)}-1+o(1)\Big{]}$ (315)
$=-C\log\Big{(}\frac{C}{k_{\nu}}\Big{)}-(1-C)\log\Big{(}\frac{1-C}{k_{0}-k_{\nu}}\Big{)}+\log\Big{(}\frac{1}{k_{0}}\Big{)}+o(1)$ (316)
$=-C\log\Big{(}\frac{C}{e^{-\nu}}\Big{)}-(1-C)\log\Big{(}\frac{1-C}{1-e^{-\nu}}\Big{)}+o(1)=-D(C\|e^{-\nu})+o(1),$ (317)
where in (317) we substitute $k_{0}=(1+o(1))k$ and $k_{\nu}=(1+o(1))e^{-\nu}k$. To complete the proof, we substitute (317) into (312), and then note that the leading factor $\frac{\Theta(1)}{\sqrt{C(1-C)\Delta}}$ with fixed $C\in[\Delta^{-1},1-\Delta^{-1}]$ (this follows from $C\in(0,1)$ and hence $11\leq C\Delta\leq\Delta-$ 1) can be written as $\exp(-o(\Delta))$ since $\Delta=\frac{\nu n}{k}=\Theta(\log k)\to\infty$. ∎
@@
$\mathbb{P}\Big{(}\big{|}W^{(\mathcal{J})}-\mathbbm{E}W^{(\mathcal{J})}\big{|}\geq\varepsilon\Big{)}\leq 2\exp\Big{(}-\frac{2\varepsilon^{2}k}{\nu\ell n}\Big{)},$ (318)
@@
$\mathbbm{E}W^{(\mathcal{J})}$ $=n\Big{(}1-\exp\Big{(}-\frac{\ell\nu(1+o(1))}{k}\Big{)}\Big{)}$ (319)
=$=\begin{cases}(1-e^{-\nu\alpha}+o(1))n,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }\frac{\ell}{k}\to\alpha\in(0,1]\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \frac{n\nu\ell}{k}(1+o(1)),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }\frac{\ell}{k}\to 0.\end{cases}$ (320)
@@
$W^{(\mathcal{J})}=\mathbbm{E}W^{(\mathcal{J})}+\delta^{\prime}n\sqrt{\frac{\ell}{k}},\text{ for some }|\delta^{\prime}|<\delta.$ (321)
@@
We first prove (319). We can write $W^{(\mathcal{J})}=\sum_{i=1}^{n}\mathbbm{1}(\text{test }i\text{ contains some item from }\mathcal{J}),$ which leads to
$\mathbbm{E}W^{(\mathcal{J})}$ $=n\cdot\mathbb{P}(\text{test }i\text{ contains some item from }\mathcal{J})$ (322)
$=n\Big{[}1-\Big{(}1-\frac{1}{n}\Big{)}^{\ell\Delta}\Big{]}$ (323)
$=n\Big{[}1-\exp\Big{(}\frac{\ell\nu n}{k}\log\big{(}1-\frac{1}{n}\big{)}\Big{)}\Big{]}$ (324)
$=n\Big{[}1-\exp\Big{(}-\frac{\ell\nu(1+o(1))}{k}\Big{)}\Big{]},$ (325)
where we use $\log(1-\frac{1}{n})=-\frac{1}{n}(1+o(1))$ in (325). Thus, (319) follows, and to obtain (320) we substitute $\frac{\ell}{k}=\alpha+o(1)$ if $\frac{\ell}{k}\to\alpha\in(0,1]$, or $1-\exp(-\beta)=\beta(1+o(1))$ as $\beta\to 0$ if $\frac{\ell}{k}=o(1)$.
Next, we establish (318). Because the $\ell\Delta$ placements of items in $\mathcal{J}$ are independent and changing one placement changes the quantity $W^{(\mathcal{J})}$ by at most one, McDiarmid’s inequality [34] gives
@@
for any $\varepsilon>0$, which yields (318) by substituting $\ell\Delta=\frac{\nu\ell n}{k}$. Finally, we show (321) by using (318). To this end, we take $\varepsilon=\delta n\sqrt{\frac{\ell}{k}}$ in (318), and note that the event in (321) holds true if and only if $|W^{(\mathcal{J})}-\mathbbm{E}W^{(\mathcal{J})}|<\delta n\sqrt{\frac{\ell}{k}}$. ∎
@@
$\mathbb{P}(M=m)\leq\exp\left(-\ell\Delta\Big{[}D\Big{(}\frac{m}{\ell\Delta}\Big{\|}\frac{n_{2}}{n}\Big{)}+\frac{\ell\Delta}{4n_{1}}\Big{]}\right),$ (326)
@@
Our argument builds on [29, Sec. IV-C], which characterizes a similar distribution (for different purposes) under the near-constant weight design. Our Lemma 8 generalizes the one therein in two ways. First, we consider the placement of $\ell$ items, while [29, Lemma 4] only concerns the case that $\ell=1$. Second, our statement is more general since we leave $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ unspecified, while [29, Lemma 4] concerns a more specific choice. As was done in [29], we make use of the Stirling numbers of the second kind, defined as {${}{n}{k}:=\frac{1}{k!}\sum_{j=0}^{k}(-1)^{k-j}\binom{k}{j}j^{n}$, e.g., see [31, P. 204].
The probability of the event that exactly $u$ placements fall in $\mathcal{N}_{1}$ is $\mathbb{P}\big{(}\mathrm{Bin}(\ell\Delta,\frac{n_{1}}{n})=u\big{)}$. Conditioning on this event, we need to calculate the probability that the remaining $\ell\Delta-u$ placements are at $m$ distinct tests in $\mathcal{N}_{2}$. Evidently, we require $\ell\Delta-u\geq m$, or equivalently $u\leq\ell\Delta-m$, in order to have $M=m$. Note that there are overall $n_{2}^{\ell\Delta-u}$ possibilities for performing $\ell\Delta-u$ placements in $n_{2}$ tests. By a standard counting argument, we can choose these $m$ distinct tests with $\binom{n_{2}}{m}$ ways, then {${}{\ell\Delta-u}{m}$ ways of arranging the $\ell\Delta-u$ placements into these $m$ unlabelled bins (which represent the $m$ tests) such that none of them are empty, and finally $m!$ different labellings of the bins. These findings give
$\mathbb{P}(M=m)$ $=\sum_{u=0}^{\ell\Delta-m}\mathbb{P}\big{(}\mathrm{Bin}\big{(}\ell\Delta,\frac{n_{1}}{n}\big{)}=u\big{)}\cdot\mathbb{P}(\ell\Delta-u\text{ placements occupy}\text{ $m$ tests in }\mathcal{N}_{2})$ (327)
@@
$\leq\sum_{t=0}^{\ell\Delta-m}\binom{\ell\Delta}{m+t}\binom{t+m}{m}m^{t}n_{1}^{\ell\Delta-m-t}$ (331)
$=\binom{\ell\Delta}{m}n_{1}^{\ell\Delta-m}\sum_{t=0}^{\ell\Delta-m}\binom{\ell\Delta-m}{t}\Big{(}\frac{m}{n_{1}}\Big{)}^{t}$ (332)
$=\binom{\ell\Delta}{m}n_{1}^{\ell\Delta-m}\Big{(}1+\frac{m}{n_{1}}\Big{)}^{\ell\Delta-m}$ (333)
::: $=C\binom{\ell\Delta}{m}n_{1}^{\ell\Delta-m},$ (334)
where we use a change of the variable $t=\ell\Delta-m-u$ in (330), {${}{t+m}{m}\leq\binom{t+m}{m}m^{t}$ [36] in (331), $\binom{\ell\Delta}{m+t}\binom{t+m}{m}=\binom{\ell\Delta}{m}\binom{\ell\Delta-m}{t}$ in (332), and $\sum_{t=0}^{\ell\Delta-m}\binom{\ell\Delta-m}{t}\big{(}\frac{m}{n_{1}}\big{)}^{t}=\big{(}1+\frac{m}{n_{1}}\big{)}^{\ell\Delta-m}$ in (333); finally, in (334), $C:=(1+\frac{m}{n_{1}})^{\ell\Delta-m}$ is bounded by
$C$ $=\exp\Big{(}[\ell\Delta-m]\log\Big{(}1+\frac{m}{n_{1}}\Big{)}\Big{)}\leq\exp\Big{(}\frac{m[\ell\Delta-m]}{n_{1}}\Big{)}\leq\exp\Big{(}\frac{\ell^{2}\Delta^{2}}{4n_{1}}\Big{)}.$ (335)
Substituting (330) back into (327), we obtain
$\mathbb{P}\big{(}M=m\big{)}$ $\leq C\binom{\ell\Delta}{m}n_{1}^{\ell\Delta-m}\frac{\binom{n_{2}}{m}m!}{n^{\ell\Delta}}$ (336)
$\leq C\binom{\ell\Delta}{m}\frac{n_{1}^{\ell\Delta-m}n_{2}^{m}}{n^{\ell\Delta}}$ (337)
$=\exp\Big{(}\frac{\ell^{2}\Delta^{2}}{4n_{1}}\Big{)}\mathbb{P}\Big{(}\mathrm{Bin}\big{(}\ell\Delta,\frac{n_{2}}{n}\big{)}=m\Big{)}$ (338)
$\leq\exp\Big{(}\frac{\ell^{2}\Delta^{2}}{4n_{1}}-\ell\Delta\cdot D\Big{(}\frac{m}{\ell\Delta}\Big{\|}\frac{n_{2}}{n}\Big{)}\Big{)},$ (339)
where we use $\binom{n_{2}}{m}m!\leq n_{2}^{m}$ in (337), (335) in (338), and the Chernoff bound (see (23) in Lemma 1) in (339). This completes the proof. ∎

### 23 Oct 2024 v2

@@ Appendix G Technical Lemmas for Near-Constant Weight Designs ($-\ell$)
$\mathbb{P}\big{(}U=C\Delta\big{)}=\exp\big{(}-\Delta\big{[}D(C\|e^{-\nu})+o(1)\big{]}\big{)}.$ (G.1)
@@
$\mathbb{P}(U=C\Delta)=\frac{\binom{k_{\nu}\Delta}{C\Delta}\binom{(k_{0}-k_{\nu})\Delta}{(1-C)\Delta}}{\binom{k_{0}\Delta}{\Delta}}.$ (G.2)
Given positive integers $m$ and $\alpha m$ for some $\alpha\in(0,1)$, we have $\frac{\sqrt{\pi}}{2}G\leq\binom{m}{\alpha m}\leq G$ where $G=\frac{\exp(mH_{2}(\alpha))}{\sqrt{2\pi m\alpha(1-\alpha)}}$ (see Lemma A.2), which gives
@@
Substituting this into (G.2), we obtain
$\mathbb{P}\big{(}U=C\Delta\big{)}$ $=\Theta(1)\frac{\frac{\exp(k_{\nu}\Delta H_{2}(\frac{C}{k_{\nu}}))}{\sqrt{C\Delta(1-\frac{C}{k_{\nu}})}}\cdot\frac{\exp((k_{0}-k_{\nu})\Delta H_{2}(\frac{1-C}{k_{0}-k_{\nu}}))}{\sqrt{(1-C)\Delta(1-\frac{1-C}{k_{0}-k_{\nu}})}}}{\frac{\exp(k_{0}\Delta H_{2}(\frac{1}{k_{0}}))}{\sqrt{\Delta(1-\frac{1}{k_{0}})}}}$ (G.3)
$=\frac{\Theta(1)}{\sqrt{C(1-C)\Delta}}\exp\Big{(}\Delta\Big{[}\underbrace{k_{\nu}H_{2}\Big{(}\frac{C}{k_{\nu}}\Big{)}+(k_{0}-k_{\nu})H_{2}\Big{(}\frac{1-C}{k_{0}-k_{\nu}}\Big{)}-k_{0}H_{2}\Big{(}\frac{1}{k_{0}}\Big{)}}_{:=\mathcal{T}}\Big{]}\Big{)},$ (G.4)
where (G.4) follows from $1-\frac{C}{k_{\nu}}$, $1-\frac{1-C}{k_{0}-k_{\nu}}$, and $1-\frac{1}{k_{0}}$ all behaving as $\Theta(1)$. Now we seek to further simplify $\mathcal{T}$. Note that $\frac{C}{k_{\nu}}$, $\frac{1-C}{k_{0}-k_{\nu}}$, and $\frac{1}{k_{0}}$ all scale as $o(1)$, and it holds for any $t=o(1)$ that
$H_{2}(t)$ $=-t\log t-(1-t)\log(1-t)=-t(\log t-1+o(1)).$ (G.5)
@@
$\mathcal{T}$ $=k_{\nu}\Big{(}-\frac{C}{k_{\nu}}\Big{[}\log\Big{(}\frac{C}{k_{\nu}}\Big{)}-1+o(1)\Big{]}\Big{)}-(k_{0}-k_{\nu})\cdot\frac{1-C}{k_{0}-k_{\nu}}\cdot\Big{[}\log\Big{(}\frac{1-C}{k_{0}-k_{\nu}}\Big{)}-1+o(1)\Big{]}$ (G.6)
$\quad\quad\quad\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +k_{0}\cdot\frac{1}{k_{0}}\cdot\Big{[}\log\Big{(}\frac{1}{k_{0}}\Big{)}-1+o(1)\Big{]}$ (G.7)
$=-C\log\Big{(}\frac{C}{k_{\nu}}\Big{)}-(1-C)\log\Big{(}\frac{1-C}{k_{0}-k_{\nu}}\Big{)}+\log\Big{(}\frac{1}{k_{0}}\Big{)}+o(1)$ (G.8)
$=-C\log\Big{(}\frac{C}{e^{-\nu}}\Big{)}-(1-C)\log\Big{(}\frac{1-C}{1-e^{-\nu}}\Big{)}+o(1)$ (G.9)
$=-D(C\|e^{-\nu})+o(1),$ (G.10)
where in (G.9) we substitute $k_{0}=(1+o(1))k$ and $k_{\nu}=(1+o(1))e^{-\nu}k$. To complete the proof, we substitute (G.10) into (G.4), and then note that the leading factor $\frac{\Theta(1)}{\sqrt{C(1-C)\Delta}}$ with fixed $C\in[\Delta^{-1},1-\Delta^{-1}]$ (this follows from $C\in(0,1)$ and hence $11\leq C\Delta\leq\Delta-$ 1) can be written as $\exp(-o(\Delta))$ since $\Delta=\frac{\nu n}{k}=\Theta(\log k)\to\infty$. ∎
@@
$\mathbb{P}\Big{(}\big{|}W^{(\mathcal{J})}-\mathbbm{E}W^{(\mathcal{J})}\big{|}\geq\varepsilon\Big{)}\leq 2\exp\Big{(}-\frac{2\varepsilon^{2}k}{\nu\ell n}\Big{)},$ (G.11)
@@
$\mathbbm{E}W^{(\mathcal{J})}$ $=n\Big{(}1-\exp\Big{(}-\frac{\ell\nu(1+o(1))}{k}\Big{)}\Big{)}$ (G.12)
=$=\begin{cases}(1-e^{-\nu\alpha}+o(1))n,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }\frac{\ell}{k}\to\alpha\in(0,1]\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \frac{n\nu\ell}{k}(1+o(1)),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }\frac{\ell}{k}\to 0.\end{cases}$ (G.13)
@@
$W^{(\mathcal{J})}=\mathbbm{E}W^{(\mathcal{J})}+\delta^{\prime}n\sqrt{\frac{\ell}{k}},\text{ for some }|\delta^{\prime}|<\delta.$ (G.14)
@@
We first prove (G.12). We can write $W^{(\mathcal{J})}=\sum_{i=1}^{n}\mathbbm{1}(\text{test }i\text{ contains some item from }\mathcal{J}),$ which leads to
$\mathbbm{E}W^{(\mathcal{J})}$ $=n\cdot\mathbb{P}(\text{test }i\text{ contains some item from }\mathcal{J})$ (G.15)
$=n\Big{[}1-\Big{(}1-\frac{1}{n}\Big{)}^{\ell\Delta}\Big{]}$ (G.16)
$=n\Big{[}1-\exp\Big{(}\frac{\ell\nu n}{k}\log\big{(}1-\frac{1}{n}\big{)}\Big{)}\Big{]}$ (G.17)
$=n\Big{[}1-\exp\Big{(}-\frac{\ell\nu(1+o(1))}{k}\Big{)}\Big{]},$ (G.18)
where we use $\log(1-\frac{1}{n})=-\frac{1}{n}(1+o(1))$ in (G.18). Thus, (G.12) follows, and to obtain (G.13) we substitute $\frac{\ell}{k}=\alpha+o(1)$ if $\frac{\ell}{k}\to\alpha\in(0,1]$, or $1-\exp(-\beta)=\beta(1+o(1))$ as $\beta\to 0$ if $\frac{\ell}{k}=o(1)$.
Next, we establish (G.11). Because the $\ell\Delta$ placements of items in $\mathcal{J}$ are independent and changing one placement changes the quantity $W^{(\mathcal{J})}$ by at most one, McDiarmid’s inequality [34] gives
@@
for any $\varepsilon>0$, which yields (G.11) by substituting $\ell\Delta=\frac{\nu\ell n}{k}$. Finally, we show (G.14) by using (G.11). To this end, we take $\varepsilon=\delta n\sqrt{\frac{\ell}{k}}$ in (G.11), and note that the event in (G.14) holds true if and only if $|W^{(\mathcal{J})}-\mathbbm{E}W^{(\mathcal{J})}|<\delta n\sqrt{\frac{\ell}{k}}$. ∎
@@
$\mathbb{P}(M=m)\leq\exp\left(-\ell\Delta\Big{[}D\Big{(}\frac{m}{\ell\Delta}\Big{\|}\frac{n_{2}}{n}\Big{)}+\frac{\ell\Delta}{4n_{1}}\Big{]}\right),$ (G.19)
@@
Our argument builds on [29, Sec. IV-C], which characterizes a similar distribution (for different purposes) under the near-constant weight design. Our Lemma G.4 generalizes the one therein in two ways. First, we consider the placement of $\ell$ items, while [29, Lemma 4] only concerns the case that $\ell=1$. Second, our statement is more general since we leave $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ unspecified, while [29, Lemma 4] concerns a more specific choice. As was done in [29], we make use of the Stirling numbers of the second kind, defined as {${}{n}{k}:=\frac{1}{k!}\sum_{j=0}^{k}(-1)^{k-j}\binom{k}{j}j^{n}$, e.g., see [31, P. 204].
The probability of the event that exactly $u$ placements fall in $\mathcal{N}_{1}$ is $\mathbb{P}\big{(}\mathrm{Bin}(\ell\Delta,\frac{n_{1}}{n})=u\big{)}$. Conditioning on this event, we need to calculate the probability that the remaining $\ell\Delta-u$ placements are at $m$ distinct tests in $\mathcal{N}_{2}$. Evidently, we require $\ell\Delta-u\geq m$, or equivalently $u\leq\ell\Delta-m$, in order to have $M=m$. Note that there are overall $n_{2}^{\ell\Delta-u}$ possibilities for performing $\ell\Delta-u$ placements in $n_{2}$ tests. By a standard counting argument, we can choose these $m$ distinct tests with $\binom{n_{2}}{m}$ ways, then {${}{\ell\Delta-u}{m}$ ways of arranging the $\ell\Delta-u$ placements into these $m$ unlabeled bins (which represent the $m$ tests) such that none of them are empty, and finally $m!$ different labelings of the bins. These findings give
$\mathbb{P}(M=m)$ $=\sum_{u=0}^{\ell\Delta-m}\mathbb{P}\big{(}\mathrm{Bin}\big{(}\ell\Delta,\frac{n_{1}}{n}\big{)}=u\big{)}\cdot\mathbb{P}(\ell\Delta-u\text{ placements occupy}\text{ $m$ tests in }\mathcal{N}_{2})$ (G.20)
@@
$\leq\sum_{t=0}^{\ell\Delta-m}\binom{\ell\Delta}{m+t}\binom{t+m}{m}m^{t}n_{1}^{\ell\Delta-m-t}$ (G.24)
$=\binom{\ell\Delta}{m}n_{1}^{\ell\Delta-m}\sum_{t=0}^{\ell\Delta-m}\binom{\ell\Delta-m}{t}\Big{(}\frac{m}{n_{1}}\Big{)}^{t}$ (G.25)
$=\binom{\ell\Delta}{m}n_{1}^{\ell\Delta-m}\Big{(}1+\frac{m}{n_{1}}\Big{)}^{\ell\Delta-m}$ (G.26)
$:=C\binom{\ell\Delta}{m}n_{1}^{\ell\Delta-m},$ (G.27)
where (G.23) follows from a change of the variable $t=\ell\Delta-m-u$, (G.24) follows from {${}{t+m}{m}\leq\binom{t+m}{m}m^{t}$ [36], (G.25) follows from $\binom{\ell\Delta}{m+t}\binom{t+m}{m}=\binom{\ell\Delta}{m}\binom{\ell\Delta-m}{t}$, (G.26) follows from $\sum_{t=0}^{\ell\Delta-m}\binom{\ell\Delta-m}{t}\big{(}\frac{m}{n_{1}}\big{)}^{t}=\big{(}1+\frac{m}{n_{1}}\big{)}^{\ell\Delta-m}$, and in (G.27), $C:=(1+\frac{m}{n_{1}})^{\ell\Delta-m}$ is bounded by
$C$ $=\exp\Big{(}[\ell\Delta-m]\log\Big{(}1+\frac{m}{n_{1}}\Big{)}\Big{)}\leq\exp\Big{(}\frac{m[\ell\Delta-m]}{n_{1}}\Big{)}\leq\exp\Big{(}\frac{\ell^{2}\Delta^{2}}{4n_{1}}\Big{)}.$ (G.28)
Substituting (G.27) back into (G.20), we obtain
$\mathbb{P}\big{(}M=m\big{)}$ $\leq C\binom{\ell\Delta}{m}n_{1}^{\ell\Delta-m}\frac{\binom{n_{2}}{m}m!}{n^{\ell\Delta}}$ (G.29)
$\leq C\binom{\ell\Delta}{m}\frac{n_{1}^{\ell\Delta-m}n_{2}^{m}}{n^{\ell\Delta}}$ (G.30)
$=\exp\Big{(}\frac{\ell^{2}\Delta^{2}}{4n_{1}}\Big{)}\mathbb{P}\Big{(}\mathrm{Bin}\big{(}\ell\Delta,\frac{n_{2}}{n}\big{)}=m\Big{)}$ (G.31)
$\leq\exp\Big{(}\frac{\ell^{2}\Delta^{2}}{4n_{1}}-\ell\Delta\cdot D\Big{(}\frac{m}{\ell\Delta}\Big{\|}\frac{n_{2}}{n}\Big{)}\Big{)},$ (G.32)
where (G.30) follows from $\binom{n_{2}}{m}m!\leq n_{2}^{m}$, (G.31) follows from (G.28), and in (G.32) we apply the Chernoff bound (see (A.1) in Lemma A.1). This completes the proof. ∎